A researcher claims that 10 year olds watch 6.6 hours of TV daily with SD = 2.5 hours. You try to verify this with the following sample data of 100 and a sample mean of 6.1 hours. Test the claim of the researcher. Test at α = .01




1
Expert's answer
2022-05-11T09:41:51-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=6.6"

"H_1:\mu\not=6.6"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.01," and the critical value for a two-tailed test is "z_c = 2.5758."

The rejection region for this two-tailed test is "R = \{z: |z| > 2.5758\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{6.1-6.6}{2.5/\sqrt{100}}=-2"

Since it is observed that "|z| = 2<2.5758=z_c ," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach: The p-value is "p=2P(Z<-2)=0.0455," and since "p = 0.0455 > 0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is different than 6.6, at the "\alpha = 0.01" significance level.